OpenStudy (anonymous):

sharon is 54 inches tall. a tree in her backyard is 5 times as tall as she is. the floor of her treehouse is at a height that is twice as talll as she is. what is the differences, in inches, between the top of the tree and the floor of the treehouse?

OpenStudy (anonymous):

Let Sharon height be x.

OpenStudy (anonymous):

Express height of tree and floor of the tree hse

OpenStudy (anonymous):

in terms of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do u get?

OpenStudy (anonymous):

u still not explaining it to me

OpenStudy (anonymous):

Lowering the temperature _____ most chemical reactions. (1 point) slows down speeds up does not affect stops HELP ME

OpenStudy (anonymous):

If the tree is 5 times the height of sharon and sharon is x inches what will that make the tree height

OpenStudy (anonymous):

Locust the answer is slows down most chemical rexn

OpenStudy (anonymous):

i am trying to lead you to the answer

OpenStudy (anonymous):

Sharon = 54 Tree = 54 * 5 = 270 Floor of Tree house = 108 270 - 108 = 162 inches

OpenStudy (anonymous):

thank u pathosdebater

OpenStudy (anonymous):

This is what i was leading you to: Let Sharon height be x. tree height is 5x. floor of trees house is 2x top of tree-floor of tree house= 5x-2x=3x but x=54 inches therefore inches 3*54

OpenStudy (anonymous):

oh wanted u break it down i had the same answer but i wanted to see if i had it right. n sumbody explaining it to me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank u moyo30

OpenStudy (anonymous):

anytime